0.78 g of sodium hydroxide (NaOH) are added to a 1 L solution of water at a temperature
of 21°C in a tube separated by a semi-permeable membrane. Calculate the osmotic pressure
of the solution
1
Expert's answer
2020-04-23T13:06:44-0400

n (NaOH) = 0.78 / 40 = 0,0195 mol

C = 0,0195 M

T = 21 + 273.15 = 296.15 K


NaOH = Na+ + OH- (2 ions, i = 2)

П = 2 * 0,0195 * 0.08206 * 296.15

П = 0,95 atm